You are at a dealer's shop to buy a new sports car for $55,000. The dealer offers you a 6 percent APR, 5-year payment plan that requires you to make equal monthly payment starting from today. What will your monthly payment be?

A. $1047.90

B. $1,053.87

C. $1,058

D. $1,063.30

E. $1,072.11
Answer #1

| PV of annuity due : 55,000.00 r:f1.xr r= 0.06/12 = 0.005 镧) n=5x12= pv of annuity 1- (1+0.005) 4-59 P = 55,000/ (1+50.98419